OpenStudy (anonymous):

-2 (y - 3) = 2y - 6 one solution, no solution, or infinite solutions?

OpenStudy (anonymous):

One solution

OpenStudy (anonymous):

y = 3

OpenStudy (arbershabani97):

-2y+6=2y-6 -4y=-6-6 y=-12/(-4) y=3 so only that solution works, if there is 3=3 infinite, if there is 4=3 then no solutions